OpenStudy (anonymous):

Hiram raises earthworms. In a square of compost 4 ft by 4 ft, he can have 1,000 earthworms. How many earthworms can he have if his square of compost has a side length that is 7 times longer? Help please?

OpenStudy (anonymous):

hiram has a lot of worms 4 x 4 = 1,000 4 x 4(7) =7 boxes at 4x4 or 7,000 worms

OpenStudy (anonymous):

above, 4 x 4 means 4 by 4, not 4 * 4, that looks confusing.

OpenStudy (anonymous):

answer choices are 49,000 or 196,000

OpenStudy (anonymous):

if the side length become 7 times ,then it should be 4(7) x 4(7)

OpenStudy (anonymous):

=49(1000)=49000

OpenStudy (anonymous):

if both side lengths are increased by 7, he would have 7*7 more worms, or 49,00

OpenStudy (anonymous):

im confused :(

OpenStudy (anonymous):

no of earthworm is proportional to the area of the plot 1000 ~ 4*4 (let ~ stand for proportional) now sides are 7 times longer (7*4)*(7*4) =7*7*(4*4) we already know (4*4)~1000 so =7*7*1000=49000

OpenStudy (anonymous):

did u get that ?

OpenStudy (anonymous):

yes thanx!:) so much

OpenStudy (anonymous):

:)